0 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

This problem is based on the concept of pH and Le Chatelier's principle which states that “when the reaction is already there at equilibrium and it is subjected to a change of concentration, pressure or temperature, the equilibrium shifts in the direction that tends to undo the effect of the change”.

Fundamentals

Le Chatelier's principle states that “when the reaction is already there at equilibrium and it is subjected to a change of concentration, pressure or temperature, the equilibrium shifts in the direction that tends to undo the effect of the change”

pH is defined as the measure of the hydrogen ion concentration of the solution.

It is calculated as follows:

pH=log[H+]{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right]

Ionization of weak acid is as follows:

On adding KNO2{\rm{KN}}{{\rm{O}}_2}, it will furnish two ions K+{{\rm{K}}^ + } and NO2{\rm{NO}}_2^ - . Therefore, when KNO2{\rm{KN}}{{\rm{O}}_2} is dissolved in the solution the NO2{\rm{NO}}_2^ - concentration increases.

So, when KNO2{\rm{KN}}{{\rm{O}}_2} is added, equilibrium shifts to left according to Le Chatelier's principle.

As the equilibrium shifts towards left, the extent of HNO2{\rm{HN}}{{\rm{O}}_2} ionization decreases.

Since the equilibrium shifts towards left, the extent of HNO2{\rm{HN}}{{\rm{O}}_2} ionization decreases. This leads to a decrease in concentration of H+{{\rm{H}}^ + }.

Since, pH=log[H+]{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right], when H+{{\rm{H}}^ + } concentration decreases, the pH value increases.

Ans:

The extent of HNO2{\rm{HN}}{{\rm{O}}_2} ionization decreases.

The concentration of H+{{\rm{H}}^ + } decreases.

The pH value increases.